ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 214, 2016

Abedelfatah, A. Hilbert functions of monomial ideals containing a regular sequence

857

Adams, T. M. and C. E. Silva Weak rational ergodicity does not imply rational ergodicity

491

Alon, N., A. Kostochka, B. Reiniger, D. B. West and X. Zhu Coloring, sparseness and girth

315

Ara, P. and E. Pardo Primely generated refinement monoids  379

Arregi, J. M. See A. Vera-López, Asensio, P. A. G., S. Sahinkaya and A. K. Srivastava New char-

acterizations of pseudo-Frobenius rings and a generalization of the FGF conjecture

121

Bae, S. and H. Jung The statistics of the number of rational points in the hyperelliptic ensemble over a finite field

421

Balister, P., B. Bollobás, Z. Füredi, I. Leader and M. Walters Subtended angles

995

Bargetz, C. and M. Dymond σ-porosity of the set of strict contractions in a space of non-expansive mappings

235

Basova, O. See A. Vera-López Bate, M., B. Martin and G. Röhrle On a question of Külshammer

for representations of finite groups in reductive groups

463

Bayart, F., R. Ernst and Q. Menet Non-existence of frequently hypercyclic subspaces for P (D)

149

Beanland, K., R. Causey and P. Motakis Arbitrarily distortable Banach spaces of higher order

553

Ben-Neria, O. The structure of the Mitchell order—I  945

Bergelson, V. and D. Robertson Polynomial recurrence with large intersection over countable fields

109

Bernig, A. and F. Voide Spin-invariant valuations on the octonionic plane

831

Bögelein, V., F. Duzaar and U. Gianazza Sharp boundedness and continuity results for the singular porous medium equation

259

Bohata, M., J. Hamhalter and O. F. K. Kalenda On Markushevich bases in preduals of von Neumann algebras

867

Bohman, T., D. Mubayi and M. Picollelli The independent neighborhoods process 333 Bollobás, B. See P. Balister Bou-Rabee, K. and D. Studenmund Full residual finiteness

growths of nilpotent groups

209

Bridson, M. R., M. Conder and A. Reid Determining Fuchsian groups by their finite quotients

1

Burton, B. and J. Spreer Combinatorial Seifert fibred spaces with transitive cyclic automorphism group

741

Causey, R. See K. Beanland Cirilo, P., Y. Lima and E. Pujals Ergodic properties of skew prod-

ucts in infinite measure

43

Conder, M. See M. R. Bridson Durango, F., J. L. Fernández and P. Fernández A limit theo-

rem for continuous selectors

983

Duzaar, F. See V. Bögelein Dymond, M. See C. Bargetz Ernst, R. See F. Bayart Farkas, Á. Projections of self-similar sets with no separation condi-

tion

67

Fehm, A. and F. Jahnke Fields with almost small absolute Galois group

193

Fernández, J. L. See F. Durango Fernández, P. See F. Durango Filmus, Y., H. Hatami, N. Keller and N. Lifshitz On the sum

of the L_{1} influences of bounded functions

167

Finster, M. Congruence Veech groups  885

Füredi, Z. See P. Balister Gianazza, U. See V. Bögelein Hamhalter, J. See M. Bohata Hatami, H. See Y. Filmus Jahnke, F. See A. Fehm Jung, H. See S. Bae Kalenda, O. F. K. See M. Bohata Keller, N. See Y. Filmus

Koenigsmann, J. On a question of Abraham Robinson  931

Kostochka, A. See N. Alon Krupski, M. and W. Marciszewski A metrizable X with C_{p}(X)

not homeomorphic to C_{p}(X) × C_{p}(X)

245

Leader, I. See P. Balister

Legrand, F. Specialization results and ramification conditions  621

Lifshitz, N. See Y. Filmus Lima, Y. See P. Cirilo Marciszewski, W. See M. Krupski Martin, B. See M. Bate Medini, A., J. van Mill and L. Zdomskyy A homogeneous space

whose complement is rigid

583

Menet, Q. See F. Bayart van Mill, J. See A. Medini Motakis, P. See K. Beanland Mubayi, D. See T. Bohman Ormaetxea, L. See A. Vera-López, van Oystaeyen, F. See X. Yu Pardo, E. See P. Ara Peralta-Salas, D., Á del Pino and F. Presas Foliated vector

fields without periodic orbits

443

Peres, Y. and M. Ṙams Projections of the natural measure for percolation fractals

539

Picollelli, M. See T. Bohman del Pino, Á. See D. Peralta-Salas Presas, F. See D. Peralta-Salas Pujals, E. See P. Cirilo Rams, M. ˙ See Y. Peres Reid, A. See M. R. Bridson Reiniger, B. See N. Alon Robertson, D. See V. Bergelson Röhrle, G. See M. Bate Sahinkaya, S. See P. A. G. Asensio

Shelah, S. Beginning of stability theory for polish spaces  507

Shu, B. See Y. Zeng Silva, C. E. See T. M. Adams Spreer, J. See B. Burton Srivastava, A. K. See P. A. G. Asensio Stern, A. and G. Urzúa KSBA surfaces with elliptic quotient sin-

gularities, π_{1} = 1, p_{g} = 0, and K^{2} = 1, 2

651

Studenmund, D. See K. Bou-Rabee Urzúa, G. See A. Stern Vera-López, A., J. M. Arregi, L. Ormaetxea, F. J. Vera-López

and O. Basova Equalities on the number of conjugacy classes of the Sylow p-subgroups of GL(n, q)

675

Vera-López, F. J. See A. Vera-López, Voide, F. See A. Bernig Walters, M. See P. Balister West, D. B. See N. Alon Wu, C. Deloné property of the holonomy vectors of translation sur-

faces

733

Yang, Y. Ramanujan-type identities for Shimura curves  699
Yoo, H. On Eisenstein ideals and the cuspidal group of J_{0}(N)  359

Yu, X., F. V. Oystaeyen and Y. Zhang Cleft extensions of Koszul twisted Calabi–Yau algebras

785

Zdomskyy, L. See A. Medini Zeng, Y. and B. Shu On Kac–Weisfeiler modules for general and

special linear Lie superalgebras

471

Zhang, Y. See X. Yu

Zhu, J. Dirichlet problem of quaternionic Monge–Ampère equations  597

Zhu, X. See N. Alon